# YOLO-LLTS: Real-Time Low-Light Traffic Sign Detection via Prior-Guided Enhancement and Multi-Branch Feature Interaction

### v3

@@ Abstract
Traffic sign detection is essential for autonomous driving and Advanced Driver Assistance Systems (ADAS). However, existing methods struggle with low-light conditions due to issues like indistinct small-object features, limited feature interaction, and poor image quality, which degrade detection accuracy and speed. To address this issue, we propose YOLO-LLTS, an end-to-end real-time traffic sign detection algorithm specifically designed for low-light environments. YOLO-LLTS introduces three main contributions: the High-Resolution Feature Map for Small Object Detection (HRFM-SOD) module to enhance small-object detection by mitigating feature dilution; the Multi-branch Feature Interaction Attention (MFIA) module to improve information extraction through multi-scale features interaction; and the Prior-Guided Feature Enhancement Module (PGFE) to enhance image quality by addressing noise, low contrast, and blurriness. Additionally, we construct a novel dataset, the Chinese Nighttime Traffic Sign Sample Set (CNTSSS), covering diverse nighttime scenarios. Experiments show that YOLO-LLTS achieves state-of-the-art performance, outperforming previous best methods by 2.7% mAP50 and 1.6% mAP50:95 on TT100K-night, 1.3% mAP50 and 1.9% mAP50:95 on CNTSSS, 7.5% mAP50 and 9.8% mAP50:95 on GTSDB-night, and superior results on CCTSDB2021. Deployment on edge devices confirms its real-time applicability and effectiveness.
Index Terms:
Traffic Sign Detection, Low-Light Conditions, Traffic Sign Dataset, End-to-End Algorithm, Edge Device Deployment.I Introduction
@@
To address the dual challenges of small targets with low resolution and low visibility in low-light environments, we have designed an end-to-end traffic sign recognition algorithm specifically for low-light conditions. As shown in Fig. 2, low-light conditions can impair the driver’s visibility, leading to the neglect of traffic signs and potential violations of traffic regulations. The ADAS, equipped with our YOLO-LLTS model, alerts the driver to take corrective actions or directly controls the vehicle upon detecting traffic signs. For example, when the system detects a speed limit sign, it automatically adjusts the vehicle’s speed to comply with the road’s speed limit. Similarly, if a stop sign is detected, the system either alerts the driver or automatically applies the brakes. YOLO-LLTS improved YOLOv8 by utilizing high-resolution images to extract clearer features, and it also designed a new multi-branch features interaction attention module to fuse features from different receptive fields. Additionally, we have developed a prior-providing module that not only enhances the image but also supplements its details. Our algorithm effectively mitigates the poor performance of existing methods under low-light conditions, thereby improving the safety of autonomous driving systems. Furthermore, to address the lack of low-light scene data in existing traffic sign datasets, we have constructed a multi-scene Chinese traffic sign dataset ranging from dusk to deep night, providing a foundational resource for industry research in this area.
@@
Multi-branch Feature Interaction Attention Module (MFIA): To enhance the model’s capability of capturing information from multiple receptive fields, we introduce the MFIA module. This module facilitates deep interaction and fusion of multi-scale features across both channel and spatial dimensions. Unlike previous methods that focus solely on single-scale attention mechanisms, MFIA effectively integrates multi-scale, semantically diverse features.
@@ III Methods
In this section, we provide a detailed introduction to our dataset and modules. Firstly, we describe the foundational aspects and detailed information of the dataset, which is intended to support further research in this area. Secondly, we introduce the design of the HRFM-SOD module, which alleviates the issue of blurred features of small objects in low-light conditions. Thirdly, we explain the design of the MFIA module, which enhances the model’s ability to perform feature interactions across multiple receptive fields. Finally, we present the PGFE module, which utilizes prior knowledge to improve the quality of low-light images.
@@ III-A Chinese Nighttime Traffic Sign Sample Set
Furthermore, in order to explore the characteristics of the CNTSSS dataset, we conducted an investigation into the object anchor box sizes within the dataset. As shown in Fig. 5, each blue dot represents the width and height of a traffic sign in pixels. The red dot marks the mean value of the dataset, with coordinates (33, 19), indicating the central tendency of the sign sizes. It can be observed that in our dataset, the traffic signs occupy a relatively small number of pixels.
III-B High-Resolution Feature Map for Small Object Detection
@@
The design details of the module architecture of YOLOv8 are shown in Fig. 7. In the feature extraction process, we adjust the receptive field obtained from the backbone using the computational method proposed in [15]. As a result, the receptive field received by the HRFM-SOD module varies with the size of the target.
@@
III-C Multi-branch Features Interaction Attention Module
In low-light conditions, small targets have low resolution and contain less information, requiring the model to have a stronger ability to capture information. The attention mechanism has been extensively studied and applied to enhance the model’s ability to capture key features (e.g., SENet [49], CANet [50], CBAM [51]). However, most existing attention mechanisms focus on processing a single feature, neglecting the potential complementarity and interactions between features. Dai et al. [52] proposed a method for attentional feature fusion of local and global features, yet this method failed to address the challenge of fusing features across more than two scales. Zhao et al. [53] proposed BA-Net, which improved SENet by leveraging information from shallow convolutional layers but overlooked the spatial domain of the image. Therefore, we propose an attention mechanism called Multi-branch Features Interaction Attention Module (MFIA).
As shown in Fig. 8, MFIA consists of two components: the Multi-branch Features Channel Attention Module (MFCA) and the Multi-branch Features Spatial Attention Module (MFSA). It effectively utilizes features with semantic and scale inconsistencies, enabling efficient interaction between multiple features in both the channel and spatial dimensions. Notably, MFIA is employed four times throughout the network. The illustration shows one MFIAi module along with its corresponding output feature .
III-C1 Multi-branch Features Channel Attention Module
@@
III-C2 Multi-branch Features Spatial Attention Module
@@
III-D Prior-Guided Feature Enhancement Module
Images captured in conditions of diminished lighting frequently often suffer from poor quality, which manifests as increased noise, reduced contrast, blurred edges, and hidden information in dark areas, significantly affecting the accuracy of traffic sign detection. However, simply using existing networks that increase exposure may amplify image noise, leaving the originally low-quality image still unclear, which negatively impacts subsequent object detection tasks.
@@ III-D Prior-Guided Feature Enhancement Module
III-D1 Prior-Guided Enhancement Module
@@
III-D2 Detail Texture Recovery Module
Feature enhancement may lead to the loss of original information in low-light images, thus requiring a network capable of effectively extracting image details and textures to supplement this. Inspired by [56, 57, 58, 59], the Invertible Neural Network (INN) extracts local information that is highly correlated with the high-frequency features in the frequency domain, specifically the edges and lines within the image. INN prevents information loss by dividing the input parameters into two parts, allowing the input and output features to mutually generate each other. This can be considered as a lossless feature extraction, which is particularly suitable in this context. As shown in Fig. 11, our DTR iterates the INN m times to achieve complete extraction of detail textures. Specifically, by splitting the input feature into two parts, and , the process becomes reversible, ensuring lossless transmission of detail features while simultaneously supplementing the enhanced feature map. The formula can be expressed as:
where represents the output parameters at the -th stage (). represents the input features, while represents the output features. The DTR module supplements the features extracted from the original image with detailed features to enhance the features processed by the PGE module. This design allows the PGFE module to enhance the image without losing any details.
@@ IV-C Comparisons with the State-of-the-arts
Tab. III presents a performance comparison of different traffic sign detection models on the CCTSDB2021 dataset. Despite the fact that the nighttime data used for training account for only a small portion of the dataset, with the majority consisting of daytime traffic sign data, our model still achieves outstanding results across multiple metrics. Specifically, our model reaches an accuracy of 88.8%, surpassing the second-ranked YOLO-TS by 0.7%. In terms of recall, our model achieves 81.1%, surpassing the second-ranked YOLO-TS by 0.3%. With an F1 score of 84.8, our model outperforms YOLO-TS by 0.5, reflecting a good balance between precision and recall. Our method also excels in the mAP50 and mAP50-95 metrics, which evaluate the model’s performance across different IoU thresholds, achieving 87.8% and 57.5%, surpassing YOLO-TS by 1.8% and 0.3%. With a parameter count of 10.2M and an FPS of 93.6, our model not only exceeds existing state-of-the-art models in accuracy but also demonstrates competitive processing speed, enabling rapid and accurate detection of traffic signs in practical applications.
@@
IV-D Ablation Study
@@ IV-D1 Effectiveness of HRFM-SOD
To validate the effectiveness of HRFM-SOD, we retain only this module on top of the baseline. As shown in the table, the HRFM-SOD module improves the mAP50 from 75.1% to 77.6%, an increase of 2.5%, and the mAP50:95 from 53.3% to 55.5%, an increase of 2.2%. The FPS also increases from 75.4 to 77.0. The experiment demonstrates that the HRFM-SOD module not only enhances detection performance but also accelerates model inference speed.
@@ IV-D2 Effectiveness of PGFE
To validate the effectiveness of PGFE, we conducted corresponding ablation experiments. As shown in the table, the PGFE module improves the mAP50 from 75.1% to 78.3%, an increase of 3.2%, and the mAP50:95 from 53.3% to 55.5%, an increase of 2.2%. When PGFE and HRFM-SOD are combined, the mAP50 increases by 4.4%, and the mAP50:95 increases by 5.8%. The experiments demonstrate that the HRFM-SOD module significantly enhances low-quality images, resulting in better detection performance. Additionally, we conducted experiments with different parameter settings in the formula. As shown in the table, the model performs best when and are set to 2 and 2.5, respectively.
@@ IV-D3 Effectiveness of MFIA
Since the MFIA module can only be used in conjunction with the HRFM-SOD module, we validated the combined effect of both modules. As shown in the second to last row of the table, the mAP50 and mAP50:95 increased by 4.5% and 5.1%, respectively, compared to the baseline. When compared to the HRFM-SOD-only configuration, mAP50 and mAP50:95 increased by 2.0% and 2.9%, respectively. The experiment demonstrates that the MFIA module effectively integrates multi-receptive field features, improving detection performance.

### v4

# YOLO-LLTS: Real-Time Low-Light Traffic Sign Detection via Prior-Guided Enhancement and Multibranch Feature Interaction

@@ Abstract
Traffic sign detection is essential for autonomous driving and Advanced Driver Assistance Systems (ADAS). However, existing methods struggle to address the challenges of poor image quality and insufficient information under low-light conditions, leading to a decline in detection accuracy and affecting driving safety. To address this issue, we propose YOLO-LLTS, an end-to-end real-time traffic sign detection algorithm specifically designed for low-light environments. YOLO-LLTS introduces three main contributions: the HRFM-SOD module retains more information about distant or tiny traffic signs compared to traditional methods; the MFIA module interacts features with different receptive fields to improve information utilization; the PGFE module enhances detection accuracy by improving brightness, edges, contrast, and supplementing detail information. Additionally, we construct a new dataset, the Chinese Nighttime Traffic Sign Sample Set (CNTSSS), covering diverse nighttime scenarios. Experiments show that YOLO-LLTS achieves state-of-the-art performance, outperforming previous best methods by 2.7% mAP50 and 1.6% mAP50:95 on TT100K-night, 1.3% mAP50 and 1.9% mAP50:95 on CNTSSS, 7.5% mAP50 and 9.8% mAP50:95 on GTSDB-night, and superior results on CCTSDB2021. Deployment on edge devices confirms its real-time applicability and effectiveness. The code and the dataset are available at https://github.com/linzy88/YOLO-LLTS.
I Introduction
@@
To address the dual challenges of small targets with low resolution and low visibility in low-light environments, we have designed an end-to-end traffic sign recognition algorithm specifically for low-light conditions. As shown in Fig. 2, low-light conditions can impair the driver’s visibility, leading to the neglect of traffic signs and potential violations of traffic regulations. The ADAS, equipped with our YOLO-LLTS model, alerts the driver to take corrective actions or directly controls the vehicle upon detecting traffic signs. For example, when the system detects a speed limit sign, it automatically adjusts the vehicle’s speed to comply with the road’s speed limit. Similarly, if a stop sign is detected, the system either alerts the driver or automatically applies the brakes. YOLO-LLTS improved YOLOv8 by utilizing high-resolution images to extract clearer features, and it also designed a new Multibranch features interaction attention module to fuse features from different receptive fields. Additionally, we have developed a prior-providing module that not only enhances the image but also supplements its details. Our algorithm effectively mitigates the poor performance of existing methods under low-light conditions, thereby improving the safety of autonomous driving systems. Furthermore, to address the lack of low-light scene data in existing traffic sign datasets, we have constructed a multi-scene Chinese traffic sign dataset ranging from dusk to deep night, providing a foundational resource for industry research in this area.
@@
Multibranch Feature Interaction Attention Module (MFIA): To enhance the model’s capability of capturing information from multiple receptive fields, we introduce the MFIA module. This module facilitates deep interaction and fusion of multi-scale features across both channel and spatial dimensions. Unlike previous methods that focus solely on single-scale attention mechanisms, MFIA effectively integrates multi-scale, semantically diverse features.
@@ III Methods
In this section, we provide a detailed introduction to our dataset and modules. Firstly, we describe the foundational aspects and detailed information of the dataset, which is intended to support further research in this area. Secondly, we introduce the overall architecture of the YOLO-LLTS model. Thirdly, We describe the design of the HRFM-SOD module, which alleviates the issue of blurred features of small objects in low-light conditions. Then, we explain the design of the MFIA module, which enhances the model’s ability to perform feature interactions across multiple receptive fields. Finally, we present the PGFE module, which utilizes prior knowledge to improve the quality of low-light images.
@@ III-A Chinese Nighttime Traffic Sign Sample Set
Furthermore, in order to explore the characteristics of the CNTSSS dataset, we conducted an investigation into the object anchor box sizes within the dataset. As shown in Fig. 5, each blue dot represents the width and height of a traffic sign in pixels. The red dot marks the mean value of the dataset, with coordinates (31, 19), indicating the central tendency of the sign sizes. It can be observed that in our dataset, the traffic signs occupy a relatively small number of pixels. To quantify the lighting differences in nighttime scenes, we calculated the grayscale values of all images in the CNTSSS and divided them into intervals. Grayscale values represent the brightness intensity of each pixel in an image. In grayscale images, each pixel’s value is typically a number ranging from 0 to 255, where 0 represents black, 255 represents white, and the intermediate values represent various shades of gray. We converted the original 3-channel BGR images to single-channel grayscale images, then computed the average grayscale value for each image. Grayscale values below 100 can be considered indicative of nighttime images. As shown in Fig. 6, images with low illumination dominate the dataset, with no images corresponding to daylight. Experiments demonstrate that our dataset presents a significant challenge for artificial intelligence models.
III-B General overview
Fig. 7 illustrates the overall architecture of the YOLO-LLTS model, which is primarily divided into three components: the backbone network, the neck, and the head.
Backbone. The backbone network is responsible for extracting initial features from the input image. It captures multi-scale important features by utilizing C2f and CBS from YOLOv8. SPPF increases the receptive field, allowing the model to capture more contextual information. The structures of C2f, CBS and SPPF are shown in Fig. 8. The first layer, the PGFE module, represents a key contribution of this work. It is integrated into the backbone network to enhance the quality of low-light images, thus improving the subsequent feature extraction process.
Neck. The neck employs the HRFM-SOD module, which we developed, responsible for multi-scale feature fusion and interaction. It receives the features extracted by the backbone network and refines these features through a series of convolutions and upsampling operations. This module is crucial for enhancing the detection of small objects. The Multibranch Feature Interaction Attention (MFIA) module is embedded within HRFM-SOD to facilitate deep interaction and fusion of multi-scale features.
Head. The head follows the design of YOLOv8 and is responsible for the final object detection and classification. It integrates the refined features from the neck and applies them to predict the position and category of traffic signs within the input image.
The main contributions focus on the PGFE module in the Backbone, the introduction of HRFM-SOD as a new Neck, and the MFIA module within HRFM-SOD. These will be detailed in the subsequent sections, respectively, with a comprehensive discussion of HRFM-SOD, MFIA, and PGFE.
III-C High-Resolution Feature Map for Small Object Detection
@@
In the feature extraction process, we adjust the receptive field obtained from the backbone using the computational method proposed in [3]. As a result, the receptive field received by the HRFM-SOD module varies with the size of the target.
@@
III-D Multibranch Features Interaction Attention Module
In low-light conditions, small targets have low resolution and contain less information, requiring the model to have a stronger ability to capture information. The attention mechanism has been extensively studied and applied to enhance the model’s ability to capture key features (e.g., SENet [20], CANet [18], CBAM [55]). However, most existing attention mechanisms focus on processing a single feature, neglecting the potential complementarity and interactions between features. Dai et al. [10] proposed a method for attentional feature fusion of local and global features, yet this method failed to address the challenge of fusing features across more than two scales. Zhao et al. [67] proposed BA-Net, which improved SENet by leveraging information from shallow convolutional layers but overlooked the spatial domain of the image. Therefore, we propose an attention mechanism called Multibranch Features Interaction Attention Module (MFIA).
As shown in Fig. 9, MFIA consists of two components: the Multibranch Features Channel Attention Module (MFCA) and the Multibranch Features Spatial Attention Module (MFSA). It effectively utilizes features with semantic and scale inconsistencies, enabling efficient interaction between multiple features in both the channel and spatial dimensions. Notably, MFIA is employed four times throughout the network. The illustration shows one MFIAi module along with its corresponding output feature .
III-D1 Multibranch Features Channel Attention Module
@@
III-D2 Multibranch Features Spatial Attention Module
@@
The MFIA module enhances the model’s ability to capture important features by facilitating interactions between features of different scales (i.e., varying sizes or resolutions). It is similar to how the human eye adjusts its focus based on the distance and importance of objects when observing a scene. In this way, the model is able to focus on critical details, especially in complex images such as small and blurry traffic signs, while reducing the interference from irrelevant parts.
III-E Prior-Guided Feature Enhancement Module
Images captured in conditions of diminished lighting often suffer from poor quality, which manifests as increased noise, reduced contrast, blurred edges, and hidden information in dark areas, significantly affecting the accuracy of traffic sign detection. However, simply using existing networks that increase exposure may amplify image noise, leaving the originally low-quality image still unclear, which negatively impacts subsequent object detection tasks.
@@ III-E Prior-Guided Feature Enhancement Module
III-E1 Prior-Guided Enhancement Module
@@
III-E2 Detail Texture Recovery Module
Feature enhancement may lead to the loss of original information in low-light images, thus requiring a network capable of effectively extracting image details and textures to supplement this. Inspired by [52, 15, 11, 68], the Invertible Neural Network (INN) extracts local information that is highly correlated with the high-frequency features in the frequency domain, specifically the edges and lines within the image. INN prevents information loss by dividing the input parameters into two parts, allowing the input and output features to mutually generate each other. This can be considered as a lossless feature extraction, which is particularly suitable in this context. As shown in Fig. 12, our DTR iterates the INN m times to achieve complete extraction of detail textures. Specifically, by splitting the input feature into two parts, and , the process becomes reversible, ensuring lossless transmission of detail features while simultaneously supplementing the enhanced feature map. Therefore, and can be expressed in terms of and , as shown in the following formula:
The next stage of is obtained by concatenating and :
where represents the output parameters at the -th stage (). As shown in Fig. 12, represents the input features, while represents the output features.
The DTR module supplements the features extracted from the original image with detailed features to enhance the features processed by the PGE module. This design allows the PGFE module to enhance the image without losing any details.
@@ IV-A4 GTSDB-night
Note: The bold blue represents the optimal result, blod green represents the suboptimal value, and differences less than or equal to 0.4% are marked with a † symbol.
@@ IV-C Comparisons with the State-of-the-arts
Tab. III presents a performance comparison of different traffic sign detection models on the CCTSDB2021 dataset. Despite the fact that the nighttime data used for training account for only a small portion of the dataset, with the majority consisting of daytime traffic sign data, our model still achieves outstanding results across multiple metrics. Specifically, our model reaches an accuracy of 88.8%, surpassing the second-ranked YOLO-TS by 0.7%. In terms of recall, our model achieves 81.1%, surpassing the second-ranked YOLO-TS by 0.3%. With an F1 score of 84.8, our model outperforms YOLO-TS by 0.5, reflecting a good balance between precision and recall. Our method also excels in the mAP50 and mAP50-95 metrics, which evaluate the model’s performance across different IoU thresholds, achieving 87.8% and 57.5%, surpassing YOLO-TS by 1.8% and 0.3%. With a parameter count of 10.2M and an FPS of 93.6, our model not only exceeds existing state-of-the-art models in accuracy but also demonstrates competitive processing speed, enabling rapid and accurate detection of traffic signs in practical applications. To rigorously assess whether the improvement is statistically significant, we trained each model 5 times with different random seeds and evaluated them on CCTSDB2021. The experiments show that the true value fluctuations for core metrics (Precision, Recall, F1, mAP50, mAP50:95) are within 0.4%, indicating that the data exhibits small and stable variations. If the difference between the optimal and suboptimal results is less than or equal to 0.4%, they can be considered statistically indistinguishable. We have marked such results with a † symbol in the top right corner.
@@
Our model has been trained and tested on both Chinese and German datasets. Experiments show that our model demonstrates a certain degree of generalization across different countries and regions. The design of our model is not dependent on the specific traffic sign styles of any particular country or region. The PGFE module, based on image enhancement and feature enhancement principles, is applicable to any low-light image. The anchor size of HRFM-SOD can be dynamically adjusted according to the target size, supporting sign proportions from different countries. When deploying the model to other countries or regions, it is recommended to fine-tune or retrain with local public datasets to further eliminate biases.
@@ IV-D1 Effectiveness of HRFM-SOD
To validate the effectiveness of HRFM-SOD, we retain only this module on top of the baseline. As shown in Tab. V, the HRFM-SOD module improves the mAP50 from 75.1% to 77.6%, an increase of 2.5%, and the mAP50:95 from 53.3% to 55.5%, an increase of 2.2%. The FPS also increases from 75.4 to 77.0. The experiment demonstrates that the HRFM-SOD module not only enhances detection performance but also accelerates model inference speed.
different and . The best result is marked in blue.
| 1.0 | 1.5 | 2.0 | 2.5 | 3.0 | |
| 1.0 | 79.7 | 79.6 | 78.1 | 79.9 | 79.8 |
| 1.5 | 80.3 | 79.6 | 80.3 | 80.0 | 79.1 |
| 2.0 | 80.1 | 80.0 | 80.7 | 81.2 | 80.0 |
| 2.5 | 80.0 | 80.2 | 80.5 | 80.6 | 79.7 |
| 3.0 | 79.2 | 79.3 | 79.1 | 78.7 | 80.3 |
| 1.0 | 1.5 | 2.0 | 2.5 | 3.0 | |
| 1.0 | 58.8 | 58.5 | 57.1 | 58.7 | 58.8 |
| 1.5 | 59.2 | 58.6 | 58.5 | 58.5 | 57.5 |
| 2.0 | 58.6 | 58.7 | 59.3 | 60.1 | 58.8 |
| 2.5 | 58.7 | 58.5 | 58.9 | 58.1 | 57.6 |
| 3.0 | 58.2 | 57.7 | 57.5 | 57.6 | 58.6 |
@@ IV-D2 Effectiveness of PGFE
To validate the effectiveness of PGFE, we conducted corresponding ablation experiments. As shown in Tab. V, the PGFE module improves the mAP50 from 75.1% to 78.3%, an increase of 3.2%, and the mAP50:95 from 53.3% to 55.5%, an increase of 2.2%. When PGFE and HRFM-SOD are combined, the mAP50 increases by 4.4%, and the mAP50:95 increases by 5.8%. The experiments demonstrate that the PGFE module plays a positive role in improving the model’s detection accuracy and makes the largest contribution to accuracy improvement among several modules. Although the FPS decreased after introducing PGFE, from 100.0 in the baseline to 74.1, it still meets the requirements for autonomous driving and driver assistance systems. Therefore, PGFE achieves significant nighttime detection gains at an acceptable computational cost, making it valuable for practical deployment.
PGFE is not a traditional pixel-level image enhancement network but a feature-level enhancement module. Therefore, evaluating the enhancement performance of PGFE cannot rely on traditional image enhancement metrics. We output features with and without PGFE, and visualized the 64-channel feature map. As shown in Fig. 15, the feature map with PGFE clearly shows sharper edges, better contrast between light and dark, and some contours that were hidden in the darkness becoming visible. This experiment further proves that PGFE can improve image quality. Additionally, we conducted experiments with different parameter settings in the formula. As shown in Tab. VI and Tab. VII, the model performs best when and are set to 2 and 2.5, respectively. The experimental range of values is set empirically to be between 1 and 3.
@@ IV-D3 Effectiveness of MFIA
Since the MFIA module can only be used in conjunction with the HRFM-SOD module, we validated the combined effect of both modules. As shown in the second to last row of Tab. V, the mAP50 and mAP50:95 increased by 4.5% and 5.1%, respectively, compared to the baseline. When compared to the HRFM-SOD-only configuration, mAP50 and mAP50:95 increased by 2.0% and 2.9%, respectively. The experiment demonstrates that the MFIA module effectively integrates multi-receptive field features, improving detection performance.
IV-E Error Analysis and Model Deployment
In this study, we conducted a detailed analysis of the training errors of the traffic sign detection model across three different datasets [42]. Fig. 16 presents the evolution of key training metrics, including normalized loss, precision, recall, mAP50, and mAP50-95, with respect to the number of training epochs. The normalized loss is the sum of the box loss, classification loss, and distribution focal Loss, which we have normalized for ease of presentation. As shown in Fig. 16, the normalized loss decreases rapidly and eventually stabilizes as the number of training epochs increases across all datasets, indicating that the model exhibits good convergence on these datasets. These critical performance metrics quickly increase in the early stages of training and eventually stabilize.
At the same time, we test the detection accuracy and inference times on mobile edge devices, NVIDIA Jetson AGX Xavier and NVIDIA Jetson AGX Orin. As shown in Fig. 17 and Fig. 18, we first connect the device to the display, keyboard, and mouse, then connect the network cable and power supply. We power on the edge computing device and test the trained weights on both the Jetson AGX Xavier and Jetson AGX Orin. The experiments on Jetson AGX Xavier were conducted in an Ubuntu 20.04 operating system environment, with optimization support provided by Jetpack 5.1. The experiments on Jetson AGX Orin were conducted in an Ubuntu 18.04 operating system environment, using the PyTorch deep learning framework (version 2.1.0), with optimization support provided by Jetpack 5.1. We selected 500 nighttime images from the CCTSDB2021 dataset for testing. As shown in Tab. VIII and Tab. IX, YOLO-LLTS achieves the best performance in terms of accuracy. Specifically, our model outperforms the average of YOLOv8 to YOLOv12 by 4.6 in accuracy, 6.7 in recall, and 5.8 in F1 score. It also exceeds the average by 7.4 in mAP50 and 4.7 in mAP50:95. On edge devices, the peak memory usage for YOLO-LLTS inference on a single 640×640 image is 88 MB, which accounts for 18% of the GPU reserved pool (88/490 MB), far from reaching the hardware limit. This result indicates that the model can still maintain usability in constrained edge environments. The number of parameters on the Xavier is 10.2 million, with an inference time of 138.8 ms. On the Orin, the number of parameters is 10.2 million, with an inference time of 44.9 ms. The experiment demonstrates that our model maintains relatively competitive performance even on devices with constrained power.